OpenStudy (anonymous):

Find the volume V of the solid obtained by rotating the region bounded by the given curves about the specified line y = 3 sin x, y = 3 cos x, 0 ≤ x ≤ π/4; about y = −1

OpenStudy (anonymous):

Here's the graph

OpenStudy (anonymous):

I know I have to use washer method by using A(x) = pi(rout)^2 - pi(rin)^2 and integrate that from 0 to pi/4 to get the volume

zepdrix (zepdrix):

Hmm yah the washer method seems like the way to go :) since both functions were given in terms of x. Hmmm let's see..

OpenStudy (anonymous):

so the outer radius would be.. 3cosx+1 and the inner radius would be 3sinx+1.. right ?

zepdrix (zepdrix):

Yessss, good good c:

OpenStudy (anonymous):

so you'd go ahead and square and have the equation pi(9cos^2x + 6cosx +1) - pi(9sin^2x + 6sinx + 1) for your area of the washer right ?

zepdrix (zepdrix):

I'd probably setup the integral before squaring, but ya that seems fine.

OpenStudy (anonymous):

Okaaay so then we'd have.. \[\Pi \int\limits_{0}^{\Pi/4}(9\cos ^{2}x + 6cosx +1) - (9\sin ^{2}x + 6sinx +1) dx\] You can pull the pi out front, right ?

zepdrix (zepdrix):

factoring it out of each term? ya looks good!

OpenStudy (psymon):

Im used totheformula naturally having pi out front xD But yeah, Im so used to picking some wrong radius or mixing something up with these, so I like to watch xD

zepdrix (zepdrix):

There is a different pi button that looks better :) lol If you're typing it use pi with lowercase P, gives you \(\large \pi\)

OpenStudy (anonymous):

Okay and then i can separate it to two integrals right ? \[\pi \int\limits_{0}^{\pi/4}(9\cos ^{2}x + 6cosx +1)dx - \pi \int\limits_{0}^{\pi/4}(9\sin ^{2}x + 6sinx + 1)dx\] Yes?

zepdrix (zepdrix):

You can, but I don't think we want to do that. I think we want to rewrite our sin^2x using our `Square Identity`. Then we can have just cos^2x's and deal with all the squares at the same time.

zepdrix (zepdrix):

\[\Large \sin^2x=1-\cos^2x\]Right? :x

OpenStudy (anonymous):

Oh yeah, okay.

OpenStudy (anonymous):

so then we'd have \[\pi \int\limits_{0}^{\pi/4}(9\cos ^{2}x + 6 cosx +1) - ((9-9\cos ^{2}x) + 6sinx +1) dx\]

OpenStudy (anonymous):

Yes ?

zepdrix (zepdrix):

yes c: simplify it down a bit! heh

OpenStudy (anonymous):

So the ones cancel and we'll have.. \[\pi \int\limits_{0}^{\pi/4}(18\cos ^{2}x + 6cosx - 6sinx - 9)dx\] right?

zepdrix (zepdrix):

good good. From there, do you remember your `Half-Angle Identity`?

OpenStudy (psymon):

Power-reducing?

zepdrix (zepdrix):

yes

OpenStudy (psymon):

Lol, yeah, thought you meant that :P

zepdrix (zepdrix):

It's called Half-Angle silly D:

OpenStudy (anonymous):

cos^2x = 1/2(1 + cos(2a)) right ?

zepdrix (zepdrix):

yah c:

OpenStudy (psymon):

There ya go :P And thats soooo power-reducing D:

zepdrix (zepdrix):

I guess I think of a different formula when I think of `Power Reducing`. When you have to solve something like:\[\Large \int\limits \cos^8x\;dx\]It involves using a formula to reduce the power a bunch of times :D

OpenStudy (psymon):

\[\cos ^{2}x=\frac{ 1+\cos2x }{ 2 }\]power-reduce \[\cos \frac{ x }{ 2 }=\pm \sqrt{\frac{ 1+cosx }{ 2 }}\]half-angle :D Lol, but don't mind me, im just being weird o.o

zepdrix (zepdrix):

They're the same formula!! :O Square the second one :3 what do you get? And see how the power is doubled in both of them? pshhhhh

zepdrix (zepdrix):

not the power, i mean the angle is doubled.. blah

OpenStudy (psymon):

lol

OpenStudy (anonymous):

so the 18 will become a 9, the 9's cancel and we have \[\pi \int\limits_{0}^{\pi/4}(9\cos(2x) + 6cosx -6sinx)dx\] correct ?

zepdrix (zepdrix):

mmmm yah that looks good.

zepdrix (zepdrix):

Smooth sailing from there :3

zepdrix (zepdrix):

These types of problems can be rather annoying. Since they're so long, it's very easy to make a mistake along the way. Do you have an answer key by chance? So we can check our answer at the end.

OpenStudy (psymon):

I usually just make a mistake with choosing the correct radii x_x

OpenStudy (anonymous):

I do not, its on webassign. But I can submit it to check it once I get there.

OpenStudy (anonymous):

So the answer is 9(pi)/2 - 6(pi)(sqrt2) ?

OpenStudy (psymon):

Let's check.

OpenStudy (psymon):

Whoops, hang on xD

zepdrix (zepdrix):

\[\Large \pi\left[\left(\frac{9}{2}+6\frac{\sqrt2}{2}+6\frac{\sqrt2}{2}\right)-\left(\color{#F35633}{6}\right)\right]\]

zepdrix (zepdrix):

When you plug in your limits, you should get something like that. Oops I simplified down the first term. but whatever. The pink value is our lower limit.

OpenStudy (anonymous):

Our lower limit.. I thought it was from 0 to pi/4 ?

zepdrix (zepdrix):

\(\large \cos0=?\)

OpenStudy (anonymous):

gahhhhh, right, its 1

OpenStudy (psymon):

\[\pi(\frac{ 9 \sin2x}{ 2 }+6sinx+6cosx)\]

OpenStudy (psymon):

Crap, my 2nd part got deleted.

zepdrix (zepdrix):

aw :c

OpenStudy (psymon):

\[\pi(\frac{ 9 }{ 2 }+\frac{ 12\sqrt{2} }{ 2 }-\frac{ 12 }{ 2 })\]

OpenStudy (psymon):

>.>

OpenStudy (anonymous):

so 9(pi)/2 + 6(pi)sqrt 2) - 6(pi)

OpenStudy (psymon):

Pretty much. Whatever simplified form you like most xD

zepdrix (zepdrix):

bah thats so ugly :3 lol

OpenStudy (psymon):

Ikr? D:

OpenStudy (anonymous):

But it's right. Yay ! Thanks guys. (:

zepdrix (zepdrix):

I simplified it down to this, I think it looks a tad nicer +_+\[\Large \frac{\pi}{2}\left[12\sqrt2-3\right]\]

zepdrix (zepdrix):

Oh good job \c:/

OpenStudy (psymon):

Okay, how id have left it doesnt look as neat as that xD